Solve the following literal equations for the indicated variable. Show all work and include the restriction.
C=R*ab/x^2; solve for x

Answer by Boreal(15235)

c=R * ab/x^2. Isolate the x
cx^2=R*ab
divide by c
x^2=R*ab/c
take the square root
x=+/- sqrt( R*ab)/c) where x cannot be 0, c cannot be 0, and Rab/c >=0.